# A JWST Transit of a Jupiter Analog I. Constraints on the Oblateness of Kepler-167 e

### v1

@@ 3.4.2 Exponential
where $d_{i}$ parameterizes the total “drop” of each exposure $i$, $b_{i}$ the starting height of each exposure $i$, and $\tau$ the length scale of the exponential drop off in the trend. For small $\tau$, this reduces to a linear model, while increasing $\tau$ forces a more aggressive exponential decay. We choose to fit for $\tau$ with large priors, but enforce that the same $\tau$ is shared across all 5 exposures. This choice was driven by initial testing that revealed letting $\tau$ vary with each exposure produced very similar values for exposures 1, 2, and 5, but markedly different values for exposures 3 and 4, where it is degenerate with the planetary parameters. By using only one $\tau$, we also add only one extra free parameter to the model compared to the linear trend. Note, however, that Sec. 5 includes a discussion on leaving out exposures 1,2, and 5 to investigate whether their contributions to $\tau$ might drive the inference on $f$, but that this had little effect. Note also that this model is distinct from the exponential model used in Paper II, which fixes $\tau$ but includes an additional linear component in time.
3.4.3 Constant (GP-only)
Our final trend model is simply a flat constant for each of the 5 exposures. Looking to Figs. 1, it is clear that this is a fairly terrible choice in isolation. However, when paired with one of the Gaussian Processes described in Sec. 3.5, this essentially gives the GP full control over both short and long term trends. Since this trend model also has 5 fewer parameters than the linear model and 6 fewer than the exponential, it is in principle possible that it ends up the preferred choice by way of model simplicity.
3.4.4 Roads Not Taken
@@ 3.5.2 Matérn 3/2
3.6 None
The final GP isn’t a GP at all: instead, we simply use the mean models from Sec. 3.4 on their own as the entire trend. This allowed us to probe the GP’s influence on the correlated residuals. We did not fit the constant mean model with no GP in light of the obvious downwards trend in each exposure.
4 MLE Results
4.1 MLE Fitting Procedure
@@
Practically, though, we were interested in maximum likelihood solutions, not maximum a-posteriori solutions, we still involved a prior in the fit to simply help keep all parameters near unity until a final transformation back to physical quantities. More specifically, we fit each parameter in an unconstrained space, then used a sigmoid to map these vectors into the unit cube, then used the transformation from these cube to physical values using the prior transformation described later in Sec. 5.1. By following this procedure we allowed the optimizer to explore all valid parameter combinations while ensuring the gradients remained numerically stable and the parameter values remained within physical limits.
Each model consisted of at least the radius ratio $R_{p}/R_{s}$, impact parameter $b$, time of mid transit $T_{0}$, projected oblateness $f$, projected ellipse rotation $\theta$, and separation $a/R_{s}$. Note that for these MLE fits, we fixed the period to a fiducial value of 1071.23205 days from Chachan et al. (2022), then fit in $a/R_{s}$, rather than in pseudo stellar density $\rho_{*}$ and period $P_{\text{orb}}$, as we do in Sec 5. In addition to these parameters, each model has 2-4 limb darkening parameters, 5-11 trend parameters, and 0 or 3 GP parameters.
@@
4.2 Intra-light curve results
@@
We also note that the point-estimates found by our MLE solver occasionally fell far from the credible ranges of the posteriors described in Sec. 5. This is in part because these fits optimized the likelihood, not the posterior, which allowed them to explore regions of parameter space at the edges of the prior. The physical limb darkening fits in particular often landed on point-esimtates where the temperature and metallicity of the star fell $5\sigma$ from the prior mean. This is also in part, however, because in these relatively high-dimensional spaces there is no guarantee that the maximum likelihood or posterior mode are within the typical set. By using a gradient-based optimizer capable of climbing narrow likelihood spikes, we can land in regions with high posterior density but negligible prior volume.
@@
4.3 Inter-light curve results
@@ 4.4 MLE Takeaways
5 Nested Sampling
5.1 Priors
@@
Secondly, when modeling $f$, the oblateness of the planet, we assumed that the distribution of true oblateness was uniform between 0 and 0.5. Above this upper bound, the centrifugal forces at the planet’s equator would overpower the gravitational attraction and the planet would come unbound (Berardo & de Wit, 2022b). However, when accounting for a planet’s precession/obliquity, a planet’s projected oblateness can only by equal to or less than its true oblateness. Berardo & de Wit (2022b) show that if one assumes that planets’ rotation axes are uniformly distributed on the sphere, the median projected $f$ is 70-75% of the true $f$.
This choice to focus on true rather than projected, oblateness may or may not be a fair assumption. On the one hand, the planets within our solar system exhibit a wide range of obliquities (from Mercury at nearly 0.0°to Uranus at nearly 90°), but on the other, one could conjure dynamical arguments that planets may be preferentially spin aligned with the their orbits (such as Mercury’s spin-orbit resonance, Murray & Dermott, 1999, though this particular mechanism would not be at play for this planet). Assuming a random distribution of orientations, however, is a more conservative choice than assuming perfect orbital alignment.
@@
5.2 Fitting Algorithm
In principle, we could have opted to use a Hamiltonian MCMC routine such as the NUTS algorithm in a package like numpyro (Hoffman & Gelman, 2011; Phan et al., 2019), since our entire model (including the numerical integrations behind squishyplanet and the stellar atmosphere interpolation within asterias) is autodifferentiable. However, we opted instead to use a nested sampling routine instead, as this allowed us to compute Bayesian evidences alongside our joint posteriors. See Skilling (2004) for the inception of nested sampling, Buchner (2023) for a recent review, and Ashton et al. (2022) for an accessible primer.
The astronomical community has access to numerous nested sampling packages in 2025, including but not limited to MultiNest (Feroz & Hobson, 2008; Feroz et al., 2009, 2019), dynesty (Speagle, 2020), and UltraNest (Buchner, 2021). Though in principle any of these options could produce our desired outputs, we opted to use nautilus, a python-based nested sampling package (Lange, 2023). During sampling, nautilus trains a simple neural network to help determine regions of improved likelihood without having to call the expensive likelihood function as often; however, when computing the evidence, it always uses the true forward model, and the bias associated with constructing bounds in this way can be removed by re-running the sampler following initial exploration. We found that nautilus converged much faster than alternative algorithms, and encourage those interested in the details of the algorithm to see Lange (2023). In contrast, we opt to use MultiNest in Paper II, as there is significant literature precedent for this algorithm in the context of exomoon transit fitting (Kipping et al., 2025).
@@ 5.2 Fitting Algorithm
5.3 Nested Sampling Results
Our highest-level takeaway from the nested sampling fits is this: when considering one light curve and using Bayes factor to perform model selection, the data are fit equally well by both spherical and oblate models. For all reductions, the Bayes factor for the best-fitting oblate and spherical models always falls below 10 (the threshold for “strong evidence”, Kass & Raftery, 1995), indicating that the additional flexibility from adding extra parameters to the model neither helps achieve markedly better fits nor dramatically hurts the model via additional complexity.
@@ 5.3 Nested Sampling Results
To place final upper bound constraints on $f$, we focus just on our oblate model fits. We first reject reductions are are overruled via a Bayes factor of 10 or more by any other model fit to that reduction. This eliminated all physical/Kurucz oblate models that used the katahdin pipeline, as each of these were overruled by spherical models with quadratic limb darkening. Next, we computed the 95th percentile of the marginalized $f$ posteriors of all remaining reductions, and selected the one with the largest value. This is came from jedi_15_6 using physical/Kurucz limb darkening, and leaves us with $f<0.097$ at 95%.
@@
6 Conclusions
6.1 Implications for Rotation Period
Having placed a bound on $f$, we can use Eq. 2 to in turn place a bound on the planet’s rotation rate $P$. As discussed in Secs. 1 and 5.1, this equation references the planet’s true oblateness, $f^{*}$, while what we have here is a bound on the projected oblateness $f$. If we temporarily assume that the true oblateness is well-matched by the projected oblateness, then we have $P\geq 7.18$ hours using Eq. 2, the posterior mode of $R_{p}/R_{s}$ from the jedi_15_6 + physical LD reduction, and the fiducial value for the star’s physical size from Chachan et al. (2022). Using an alternative formulation from de Pater & Lissauer (2015), as is done in Lammers & Winn (2024), gives $P\geq 7.11$ hours when using the same $\Lambda_{2}$ as Jupiter (0.165).
Marginalizing over the uncertainty in the planet’s true orientation is challenging in this case since posteriors from different reductions have different marginal distributions for $f$. However, we can get a sense of its effects using the random samples used to generate the prior for $f$ in Fig. 11. If we take all of the randomly-generated $f$’s that fell below $f<0.097$ and investigate their corresponding $f^{*}$’s, we find the 95th percentile is $f^{*}<0.431$, corresponding to $P\geq 3.35$ hours. If we additionally enforce that the true obliquity is $45^{\circ}$, as in Lammers & Winn (2024), we find $f^{*}<0.125$ and $P\geq 6.30$ hours. These are crude bounds included for illustrative purpose more than anything, as they do not actually marginalize over any of our posteriors and are derived only by considering our 95% upper limit on $f$. However, we believe it worth restating that every reported period bound is conditioned on an assumption about the planet’s true orientation.
6.2 Challenges and Promise of Measuring Oblateness with JWST
@@
As noted in Sec. 5, different reduction/model combinations lead to a wide range of bounds on $f$: our smallest upper bound is nearly 33% smaller than our highest upper bound. Further, examination of Tables 2 and 3 reveals that lowest reduced $\chi^{2}$ of our fits, that achieved for katahdin_snr_10, a spherical planet, and quadratic limb darkening, is 1.19. This indicates that even with our flexible GPs the models tested are still not fitting these data as well as we might expect. However, with multiple exposures to contend with, it is challenging to experiment with and fit more complex trend models in a typical Bayesian workflow. The models presented above already required fitting 21-25 free parameters and push the bounds of nested sampling routines when using a relatively expensive forward model like squishyplanet. While it may be possible to explore higher-dimensional spaces using Hamiltonian Monte Carlo routines, this would require abandoning Bayesian evidence as a model comparison tool. Paper II, which deals with even higher dimensional models involving a planet and a moon, explores using profile likelihood as an alternative technique that allows for computation of Bayesian evidence in higher dimensional spaces.
@@
This research has made use of the Astrophysics Data System, funded by NASA under Cooperative Agreement 80NSSC21M00561. We thank the Yale Center for Research Computing for guidance and assistance in the computation run on the Bouchet Cluster. Resources supporting this work were provided by the NASA High-End Computing (HEC) Program through the NASA Advanced Supercomputing (NAS) Division at Ames Research Center. This work is based [in part] on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program #6491. The data can be found in MAST: http://dx.doi.org/10.17909/e50n-4y96 (catalog 10.17909/e50n-4y96). Special thanks to donors to the Cool Worlds Lab, without whom this kind of research would not be possible: Douglas Daughaday, Elena West, Tristan Zajonc, Alex de Vaal, Mark Elliott, Stephen Lee, Zachary Danielson, Chad Souter, Marcus Gillette, Jason Rockett, Tom Donkin, Andrew Schoen, Mike Hedlund, Ryan Provost, Nicholas De Haan, Emerson Garland, Queen Rd Fnd Inc., Ieuan Williams, Axel Nimmerjahn, Brian Cartmell, Guillaume Le Saint, Robin Raszka, Bas van Gaalen, Josh Alley, Drew Aron, Warren Smith, Brad Bueche, Steve Larter, Marisol Adler & Craig Frederick.

### v2

@@ III.4.2 Exponential
where $d_{i}$ parameterizes the total “drop” of each exposure $i$, $b_{i}$ the starting height of each exposure $i$, and $\tau$ the length scale of the exponential drop off in the trend. Time $t$ is normalized between 0 and 1, and the prefactor is derived by evaluating the exponential at $t=0$ and $t=1$, which produces the apparently suppressed units. For small $\tau$, this reduces to a linear model, while increasing $\tau$ forces a more aggressive exponential decay. We choose to fit for $\tau$ with large priors, but enforce that the same $\tau$ is shared across all 5 exposures. This choice was driven by initial testing that revealed letting $\tau$ vary with each exposure produced very similar values for exposures 1, 2, and 5, but markedly different values for exposures 3 and 4, where it is degenerate with the planetary parameters. By using only one $\tau$, we also add only one extra free parameter to the model compared to the linear trend. Note, however, that Sec. V includes a discussion on leaving out exposures 1,2, and 5 to investigate whether their contributions to $\tau$ might drive the inference on $f$, but that this had little effect. Note also that this model is distinct from the exponential model used in Paper II, which fixes $\tau$ but includes an additional linear component in time.
III.4.3 Constant (GP-only)
Our final trend model is simply a flat constant for each of the 5 exposures. Looking to Figs. 1, it is clear that this is a fairly terrible choice in isolation. However, when paired with one of the Gaussian Processes described in Sec. III.5, this essentially gives the GP full control over both short and long term trends. Since this trend model also has 5 fewer parameters than the linear model and 6 fewer than the exponential, it is in principle possible that it ends up the preferred choice by way of model simplicity.
III.4.4 Roads Not Taken
@@ III.5.2 Matérn 3/2
III.6 None
The final GP isn’t a GP at all: instead, we simply use the mean models from Sec. III.4 on their own as the entire trend. This allowed us to probe the GP’s influence on the correlated residuals. We did not fit the constant mean model with no GP in light of the obvious downwards trend in each exposure.
IV MLE Results
IV.1 MLE Fitting Procedure
@@
Practically, though, we were interested in maximum likelihood solutions, not maximum a-posteriori solutions, we still involved a prior in the fit to simply help keep all parameters near unity until a final transformation back to physical quantities. More specifically, we fit each parameter in an unconstrained space, then used a sigmoid to map these vectors into the unit cube, then used the transformation from these cube to physical values using the prior transformation described later in Sec. V.1. By following this procedure we allowed the optimizer to explore all valid parameter combinations while ensuring the gradients remained numerically stable and the parameter values remained within physical limits.
Each model consisted of at least the radius ratio $R_{p}/R_{s}$, impact parameter $b$, time of mid transit $T_{0}$, projected oblateness $f$, projected ellipse rotation $\theta$, and separation $a/R_{s}$. Note that for these MLE fits, we fixed the period to a fiducial value of 1071.23205 days from Chachan et al. (2022), then fit in $a/R_{s}$, rather than in pseudo stellar density $\rho_{*}$ and period $P_{\text{orb}}$, as we do in Sec V. In addition to these parameters, each model has 2-4 limb darkening parameters, 5-11 trend parameters, and 0 or 3 GP parameters.
@@
IV.2 Intra-light curve results
@@
We also note that the point-estimates found by our MLE solver occasionally fell far from the credible ranges of the posteriors described in Sec. V. This is in part because these fits optimized the likelihood, not the posterior, which allowed them to explore regions of parameter space at the edges of the prior. The physical limb darkening fits in particular often landed on point-esimtates where the temperature and metallicity of the star fell $5\sigma$ from the prior mean. This is also in part, however, because in these relatively high-dimensional spaces there is no guarantee that the maximum likelihood or posterior mode are within the typical set. By using a gradient-based optimizer capable of climbing narrow likelihood spikes, we can land in regions with high posterior density but negligible prior volume.
@@
IV.3 Inter-light curve results
@@ IV.4 MLE Takeaways
V Nested Sampling
V.1 Priors
@@
Secondly, when modeling $f$, the oblateness of the planet, we assumed that the distribution of true oblateness was uniform between 0 and 0.5. Above this upper bound, the centrifugal forces at the planet’s equator would overpower the gravitational attraction and the planet would come unbound (Berardo and de Wit, 2022a). However, when accounting for a planet’s precession/obliquity, a planet’s projected oblateness can only by equal to or less than its true oblateness. Berardo and de Wit (2022a) show that if one assumes that planets’ rotation axes are uniformly distributed on the sphere, the median projected $f$ is 70-75% of the true $f$.
This choice to focus on true rather than projected, oblateness may or may not be a fair assumption. On the one hand, the planets within our solar system exhibit a wide range of obliquities (from Mercury at nearly 0.0°to Uranus at nearly 90°), but on the other, one could conjure dynamical arguments that planets may be preferentially spin aligned with the their orbits (such as Mercury’s spin-orbit resonance, Murray and Dermott, 1999, though this particular mechanism would not be at play for this planet). Assuming a random distribution of orientations, however, is a more conservative choice than assuming perfect orbital alignment.
@@
V.2 Fitting Algorithm
In principle, we could have opted to use a Hamiltonian MCMC routine such as the NUTS algorithm in a package like numpyro (Hoffman and Gelman, 2011; Phan et al., 2019), since our entire model (including the numerical integrations behind squishyplanet and the stellar atmosphere interpolation within asterias) is autodifferentiable. However, we opted instead to use a nested sampling routine instead, as this allowed us to compute Bayesian evidences alongside our joint posteriors. See Skilling (2004) for the inception of nested sampling, Buchner (2023) for a recent review, and Ashton et al. (2022) for an accessible primer.
The astronomical community has access to numerous nested sampling packages in 2025, including but not limited to MultiNest (Feroz and Hobson, 2008; Feroz et al., 2009, 2019), dynesty (Speagle, 2020), and UltraNest (Buchner, 2021). Though in principle any of these options could produce our desired outputs, we opted to use nautilus, a python-based nested sampling package (Lange, 2023). During sampling, nautilus trains a simple neural network to help determine regions of improved likelihood without having to call the expensive likelihood function as often; however, when computing the evidence, it always uses the true forward model, and the bias associated with constructing bounds in this way can be removed by re-running the sampler following initial exploration. We found that nautilus converged much faster than alternative algorithms, and encourage those interested in the details of the algorithm to see Lange (2023). In contrast, we opt to use MultiNest in Paper II, as there is significant literature precedent for this algorithm in the context of exomoon transit fitting (Kipping et al., 2025).
@@ V.2 Fitting Algorithm
V.3 Nested Sampling Results
Our highest-level takeaway from the nested sampling fits is this: when considering one light curve and using Bayes factor to perform model selection, the data are fit equally well by both spherical and oblate models. For all reductions, the Bayes factor for the best-fitting oblate and spherical models always falls below 10 (the threshold for “strong evidence”, Kass and Raftery, 1995), indicating that the additional flexibility from adding extra parameters to the model neither helps achieve markedly better fits nor dramatically hurts the model via additional complexity.
@@ V.3 Nested Sampling Results
To place final upper bound constraints on $f$, we focus just on our oblate model fits. We first reject reductions are overruled via a Bayes factor of 10 or more by any other model fit to that reduction. This eliminated all physical/Kurucz oblate models that used the katahdin pipeline, as each of these were overruled by spherical models with quadratic limb darkening. Next, we computed the 95th percentile of the marginalized $f$ posteriors of all remaining reductions, and selected the one with the largest value. This is came from jedi_15_6 using physical/Kurucz limb darkening, and leaves us with $f<0.097$ at 95%.
@@
VI Conclusions
VI.1 Implications for Rotation Period
Having placed a bound on $f$, we can use Eq. 2 to in turn place a bound on the planet’s rotation rate $P$. As discussed in Secs. I and V.1, this equation references the planet’s true oblateness, $f^{*}$, while what we have here is a bound on the projected oblateness $f$. If we temporarily assume that the true oblateness is well-matched by the projected oblateness, then we have $P\geq 7.18$ hours using Eq. 2, the posterior mode of $R_{p}/R_{s}$ from the jedi_15_6 + physical LD reduction, and the fiducial value for the star’s physical size from Chachan et al. (2022). Using an alternative formulation from de Pater and Lissauer (2015), as is done in Lammers and Winn (2024), gives $P\geq 7.11$ hours when using the same $\Lambda_{2}$ as Jupiter (0.165).
Marginalizing over the uncertainty in the planet’s true orientation is challenging in this case since posteriors from different reductions have different marginal distributions for $f$. However, we can get a sense of its effects using the random samples used to generate the prior for $f$ in Fig. 11. If we take all of the randomly-generated $f$’s that fell below $f<0.097$ and investigate their corresponding $f^{*}$’s, we find the 95th percentile is $f^{*}<0.431$, corresponding to $P\geq 3.35$ hours. If we additionally enforce that the true obliquity is $45^{\circ}$, as in Lammers and Winn (2024), we find $f^{*}<0.125$ and $P\geq 6.30$ hours. These are crude bounds included for illustrative purpose more than anything, as they do not actually marginalize over any of our posteriors and are derived only by considering our 95% upper limit on $f$. However, we believe it worth restating that every reported period bound is conditioned on an assumption about the planet’s true orientation.
VI.2 Challenges and Promise of Measuring Oblateness with JWST
@@
As noted in Sec. V, different reduction/model combinations lead to a wide range of bounds on $f$: our smallest upper bound is nearly 33% smaller than our highest upper bound. Further, examination of Tables 2 and 3 reveals that lowest reduced $\chi^{2}$ of our fits, that achieved for katahdin_snr_10, a spherical planet, and quadratic limb darkening, is 1.19. This indicates that even with our flexible GPs the models tested are still not fitting these data as well as we might expect. However, with multiple exposures to contend with, it is challenging to experiment with and fit more complex trend models in a typical Bayesian workflow. The models presented above already required fitting 21-25 free parameters and push the bounds of nested sampling routines when using a relatively expensive forward model like squishyplanet. While it may be possible to explore higher-dimensional spaces using Hamiltonian Monte Carlo routines, this would require abandoning Bayesian evidence as a model comparison tool. Paper II, which deals with even higher dimensional models involving a planet and a moon, explores using profile likelihood as an alternative technique that allows for computation of Bayesian evidence in higher dimensional spaces.
@@
This research has made use of the Astrophysics Data System, funded by NASA under Cooperative Agreement 80NSSC21M00561. We thank the Yale Center for Research Computing for guidance and assistance in the computation run on the Bouchet Cluster. Resources supporting this work were provided by the NASA High-End Computing (HEC) Program through the NASA Advanced Supercomputing (NAS) Division at Ames Research Center. This work is based [in part] on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program #6491. The JWST data presented in this article were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute. The specific observations analyzed can be accessed via https://doi.org/10.17909/e50n-4y96 (catalog doi: 10.17909/e50n-4y96). Special thanks to donors to the Cool Worlds Lab, without whom this kind of research would not be possible: Douglas Daughaday, Elena West, Tristan Zajonc, Alex de Vaal, Mark Elliott, Stephen Lee, Zachary Danielson, Chad Souter, Marcus Gillette, Jason Rockett, Tom Donkin, Andrew Schoen, Mike Hedlund, Ryan Provost, Nicholas De Haan, Emerson Garland, Queen Rd Fnd Inc., Ieuan Williams, Axel Nimmerjahn, Brian Cartmell, Guillaume Le Saint, Robin Raszka, Bas van Gaalen, Josh Alley, Drew Aron, Warren Smith, Brad Bueche, Steve Larter, Marisol Adler & Craig Frederick.